OpenStudy (anonymous):

Find the equation of the line through the points (1,7) and (3, -1).

OpenStudy (mimi_x3):

y = -4x + 11

OpenStudy (anonymous):

Can you explain your work? I got something different.

OpenStudy (lalaly):

\[\underline{Slope}\]\[=\large{\frac{y_2-y_1}{x_2-x_1}}\]

OpenStudy (lalaly):

\[= \frac{-1-7}{3-1} = \frac{-8}{2}=-4\]

OpenStudy (mimi_x3):

m = -1-7/3-1 = -4 y = mx + b 7 = -4(1) + b 7 = - 4 + b b = 11 y = -4x + 11

OpenStudy (lalaly):

now \[y-y_1=m(x-x_1)\]\[y-7=-4(x-1)\]

OpenStudy (lalaly):

\[y-7=-4x+4\]\[y=-4x+11\]

OpenStudy (anonymous):

Thank you both!

OpenStudy (lalaly):

\[\Huge{{your}} \Huge{\underline{Welcome}}\]

OpenStudy (lalaly):

:D